# Quantifying Uncertainty: All We Need is the Bootstrap?

### v2

@@ 3.4 Absolute distance from exact CIs
Coverage by itself is not enough - we would also like the CI endpoints to match the exact. An extreme example would be to generate a very large endpoint (wide two-sided interval) with probability $\alpha$ and a very small endpoint (narrow two-sided interval) with probability $\alpha$. This would result in nominal coverage but useles CIs.
@@
Mean absolute distance from exact for one-sided CIs. For each combination, the value is normalized with two standard deviations of exact intervals. The all column is across all experiments, while the remaining results are grouped by sample size or statistical functional. The best performing method for each column is underlined. all 4 8 16 32 64 128 256 corr mean Q0.5 Q0.05 Q0.95 std B-n $\underline{0.245}$ $0.423$ $\underline{0.337}$ $\underline{0.276}$ $\underline{0.228}$ $\underline{0.187}$ $\underline{0.151}$ $\underline{0.123}$ $\underline{0.095}$ $\underline{0.140}$ $\underline{0.201}$ $\underline{0.329}$ $\underline{0.330}$ $0.244$ $B$ $0.271$ $\underline{0.392}$ $0.354$ $0.328$ $0.270$ $0.213$ $0.185$ $0.160$ $\underline{0.095}$ $0.141$ $0.262$ $0.362$ $0.383$ $\underline{0.227}$ BC $0.322$ $0.566$ $0.436$ $0.373$ $0.316$ $0.254$ $0.210$ $0.177$ $0.136$ $0.161$ $0.286$ $0.416$ $0.536$ $0.258$ SB $0.326$ $0.546$ $0.446$ $0.362$ $0.306$ $0.252$ $0.207$ $0.172$ $0.125$ $0.148$ $0.276$ $0.479$ $0.441$ $0.308$ BCa $0.327$ $0.542$ $0.427$ $0.379$ $0.326$ $0.274$ $0.224$ $0.186$ $0.136$ $0.180$ $0.286$ $0.432$ $0.521$ $0.268$ DB $0.353$ $0.557$ $0.488$ $0.387$ $0.331$ $0.290$ $0.240$ $0.193$ $0.159$ $0.224$ $0.299$ $0.488$ $0.501$ $0.279$ PB $0.359$ $0.644$ $0.508$ $0.399$ $0.324$ $0.258$ $0.213$ $0.178$ $0.124$ $0.150$ $0.295$ $0.554$ $0.502$ $0.320$ B-t $1.018$ $4.647$ $0.702$ $0.700$ $0.550$ $0.314$ $0.242$ $0.197$ $0.116$ $0.776$ $1.751$ $0.912$ $1.127$ $0.636$
Table 3.4 shows where a baseline method has better distance from exact than DB, but only for cases where its coverage is not an order of magnitude worse. We included B-n as the bootstrap method that performs best in distance from exact. Results are similar to those in Table 3.3 - baseline methods and B-n outperform DB for the two extreme percentiles. That is, in most cases, lower distance from exact comes at the expense of worse coverage.
@@
Cases where baseline (with B-n) outperforms DB in absolute distance from exact intervals. The number in the parentheses is the number of combinations where a method outperforms. Note that there are a total of 36 combinations for every pair of $n$ and functional, except Pearson correlation, where there are 6. n functional baseline $\gg$ DB 4 corr fisher (6) 4 mean B-n (9); t-test (9); wilcoxon (9) 4 median B-n (5); m-j (5); q-nonpar (4); q-par (5); wilcoxon (5) 4 Q(0.05) B-n (27); q-nonpar (27); q-par (24) 4 Q(0.95) B-n (21); q-nonpar (10); q-par (21) 4 std B-n (14); chi-sq (14) 8 corr fisher (6) 8 mean B-n (3); t-test (3); wilcoxon (3) 8 median B-n (1); m-j (1); q-nonpar (1); q-par (1); wilcoxon (1) 8 Q(0.05) B-n (30); q-nonpar (28); q-par (24) 8 Q(0.95) B-n (27); q-nonpar (15); q-par (23) 8 std B-n (7); chi-sq (7) 16 Q(0.05) B-n (35); q-nonpar (35); q-par (27) 16 Q(0.95) B-n (16); m-j (16); q-nonpar (2); q-par (16) 16 std B-n (2); chi-sq (2) 32 Q(0.05) B-n (29); m-j (29); q-nonpar (26); q-par (15) 32 Q(0.95) B-n (30); m-j (30); q-nonpar (13); q-par (20) 64 Q(0.05) B-n (8); m-j (8); q-nonpar (8); q-par (4) 64 std B-n (1); chi-sq (1) 128 Q(0.05) B-n (2); m-j (2); q-nonpar (2); q-par (1)
@@
Figure 2: A visualization of the visualization tool for browsing the results of our simulation study. The user can filter results by method and facet based on the dimensions of the experiment (functional, confidence level, distribution).
@@ 4 Conclusion
This paper presents the most comprehensive review of empirical results for bootstrap methods to date, along with an extensive empirical comparison. Our study encompasses not only the most widely used non-parametric bootstrap methods but also the most common techniques for quantifying uncertainty in general. Furthermore, we introduce a novel criterion for assessing confidence interval quality, which improves on existing evaluation approaches.
@@
We identify DB as the overall best method, contrary to some recommendations of BCa for general cases [2], but in line with other recommendations [49, 23] and the few empirical studies that included DB. These results are also in line with related work: PB performs relatively poorly. For small $n$ B-t performs best for the mean and BCa best for Pearson correlation, although DB performs similarly. B-t relies on an estimate of variance and can produce very long CIs. Bootstrapping can perform poorly compared do chi-squared CIs on variance for small $n$ and normal distribution, however, as we demonstrated, chi-squared performs poorly on non-normal distributions and DB is comparable or better for $n\geq 32$.
@@
DB has two weaknesses. It can perform relatively poorly when $n=4,8$ and on extreme percentiles for $n\leq 32$. The latter can be mitigated by using B-n but also raises the question if DB can be modified to deal with these cases, while still preserving most of its simplicity. This leaves us with $n=4,8$, where we have to acknowledge that a non-parametric approach is often worse than a parametric approach even when the assumptions of the parametric method are violated. However, we would argue that these are extremely small and of little relevance to most practitioners. We did not investigate to what extent bootstrap diagnostics (see [50] and [3, Ch.3.10]) could further mitigate these issues.
@@
The dimensions of our experiment can be improved. We omit categorical data with commonly used baselines such as the binomial test for proportions. We only include one DGP for Pearson correlation, which is not nearly as comprehensive as the recent study by [37], who, unfortunately, did not include DB. Other commonly used functionals could also be included, such as regression model coefficients, non-parametric correlation, and distances between distributions. However, we believe that the dimensions our simulation study were diverse enough to reveal the most important results.
We focused on independent data. Related work on hierarchical, temporal, spatial, and other dependencies is sparse, but mostly in favor of bootstrapping. And in more complex problems bootstrapping is often the only viable approach. From a pedagogical perspective, the bootstrap on dependent data is more difficult, but the most of the difficulties arise from understanding the dependencies and adjusting how to resample accordingly.
There is no substitute for thoroughly understanding the statistical task at hand, carefully choosing the most appropriate method, and correctly applying it. In practice, however, these come at a training cost that is prohibitive for most practitioners. Any simplification is welcome and we identify the bootstrap as a viable one-size-fits-all method for quantifying uncertainty, at least in the independent data case. We would argue that using a single and conceptually more simple method outweighs even a moderate drop in performance compared to methods that are most commonly taught and used today. However, we show that bootstrapping even performs as well if not better on all but the smallest sample sizes.
Arguably, the lack of adoption of the bootstrap as a one-size-fits-all method might not be due to performance issues but for a lack of software, materials, and legitimate concerns for backwards compatibility. The logical next step appears to be to develop and test a first course in applied statistics that relies solely on the bootstrap.
@@ Acknowledgment
This work was supported by the Slovenian Research and Innovation Agency under Grant P2-0442.

### v3

@@ 3.4 Absolute distance from exact CIs
Coverage by itself is not sufficient, as CI endpoints should also match the exact endpoints. An extreme example would be to generate a very large endpoint (wide two-sided interval) with probability $\alpha$ and a very small endpoint (narrow two-sided interval) with probability $\alpha$. This would result in nominal coverage but useless CIs.
@@
Mean absolute distance from exact for one-sided CIs. For each combination, the value is normalized with two standard deviations of exact intervals. The all column is across all experiments, while the remaining results are grouped by sample size or statistical functional. The best performing method for each column is underlined. all 4 8 16 32 64 128 256 corr mean Q0.5 Q0.05 Q0.95 std B-n $\underline{0.240}$ $0.423$ $\underline{0.333}$ $\underline{0.269}$ $\underline{0.221}$ $\underline{0.180}$ $\underline{0.145}$ $\underline{0.118}$ $\underline{0.095}$ $\underline{0.146}$ $\underline{0.201}$ $\underline{0.329}$ $\underline{0.330}$ $0.244$ $B$ $0.264$ $\underline{0.395}$ $0.346$ $0.317$ $0.261$ $0.206$ $0.178$ $0.153$ $\underline{0.095}$ $\underline{0.146}$ $0.262$ $0.362$ $0.383$ $\underline{0.227}$ SB $0.318$ $0.544$ $0.439$ $0.352$ $0.295$ $0.243$ $0.199$ $0.165$ $0.125$ $0.161$ $0.276$ $0.479$ $0.441$ $0.308$ BC $0.332$ $0.620$ $0.461$ $0.382$ $0.316$ $0.253$ $0.208$ $0.173$ $0.136$ $0.242$ $0.286$ $0.416$ $0.536$ $0.258$ BCa $0.340$ $0.600$ $0.458$ $0.392$ $0.329$ $0.275$ $0.222$ $0.183$ $0.136$ $0.268$ $0.286$ $0.432$ $0.521$ $0.268$ PB $0.350$ $0.642$ $0.499$ $0.386$ $0.313$ $0.249$ $0.205$ $0.170$ $0.124$ $0.166$ $0.295$ $0.554$ $0.502$ $0.320$ DB $0.360$ $0.597$ $0.506$ $0.394$ $0.329$ $0.286$ $0.234$ $0.188$ $0.159$ $0.284$ $0.299$ $0.488$ $0.501$ $0.279$ B-t $0.966$ $4.376$ $0.673$ $0.668$ $0.524$ $0.300$ $0.231$ $0.188$ $0.116$ $0.624$ $1.751$ $0.912$ $1.127$ $0.636$
Table 3.4 shows where a baseline method has better distance from exact than DB, but only for cases where its coverage is not an order of magnitude worse. B-n is included as the bootstrap method that performs best in distance from exact. Results are similar to those in Table 3.3 - baseline methods and B-n outperform DB for the two extreme percentiles. That is, in most cases, lower distance from exact comes at the expense of worse coverage.
@@
Cases where baseline (with B-n included) outperforms DB in absolute distance from exact intervals. The number in the parentheses is the number of combinations where a method outperforms. Note that there are a total of 36 combinations for every pair of $n$ and functional, except Pearson correlation, where there are 6, and the mean, where there are 48. n functional baseline and B-n $\gg$ DB 4 corr fisher (6) 4 mean a-c (2); B-n (7); c-p (1); t-test (3); wilcoxon (1) 4 median B-n (2); m-j (2); q-par (2) 4 Q(0.05) B-n (19); q-nonpar (3); q-par (17) 4 Q(0.95) B-n (17); q-par (15) 4 std B-n (6) 8 corr fisher (6) 8 mean a-c (2); B-n (6); c-p (1); t-test (8); wilcoxon (2) 8 Q(0.05) B-n (21); q-nonpar (6); q-par (18) 8 Q(0.95) B-n (19); q-nonpar (7); q-par (19) 16 corr fisher (1) 16 mean B-n (4); t-test (3) 16 Q(0.05) B-n (11); q-nonpar (3); q-par (6) 16 Q(0.95) B-n (8); m-j (6); q-par (6) 32 mean a-c (1); B-n (3); c-p (1); t-test (5) 32 Q(0.05) B-n (8); m-j (5); q-par (9) 32 Q(0.95) B-n (8); m-j (4); q-par (11) 32 std chi-sq (1) 64 mean B-n (1); t-test (1) 64 Q(0.05) B-n (3); m-j (1); q-par (1) 64 Q(0.95) B-n (1); m-j (6); q-par (1) 64 std chi-sq (2) 128 mean B-n (1); t-test (1) 128 Q(0.05) B-n (1); m-j (1); q-nonpar (1) 128 Q(0.95) B-n (2); m-j (3)
@@
Figure 2: A visualization of the visualization tool for browsing the results of the simulation study. The user can filter results by method and facet based on the dimensions of the experiment (functional, confidence level, distribution).
@@ 4 Conclusion
This paper is the most comprehensive review of empirical results for bootstrap methods to date, along with an extensive empirical comparison. The simulation study encompasses not only the most widely used non-parametric bootstrap methods but also the most common techniques for quantifying uncertainty in general. Furthermore, a novel criterion for assessing confidence interval quality is introduced, which improves on existing evaluation approaches.
@@
DB is identified as the overall best method, contrary to some recommendations of BCa for general cases [2], but in line with other recommendations [51, 23] and the few empirical studies that included DB. These results are also in line with related work: PB performs relatively poorly. For small $n$ B-t performs best for the mean and BCa best for Pearson correlation, although DB performs similarly. B-t relies on an estimate of variance and can produce very long CIs. Bootstrapping can perform poorly compared do chi-squared CIs on variance for small $n$ and normal distribution, however, as demonstrated, chi-squared performs poorly on non-normal distributions and DB is comparable or better for $n\geq 32$.
@@
DB has two weaknesses. It can perform relatively poorly when $n=4,8$ and on extreme percentiles for $n\leq 32$. The latter can be mitigated by using B-n but also raises the question if DB can be modified to deal with these cases, while still preserving most of its simplicity. This leaves us with $n=4,8$, where it has to be acknowledged that a non-parametric approach is often worse than a parametric approach even when the assumptions of the parametric method are violated. It was not investigated to what extent bootstrap diagnostics (see [52] and [3, Ch.3.10]) could further mitigate these issues.
@@
The dimensions of the experiment can be improved, as it only includes one DGP for Pearson correlation, which is not nearly as comprehensive as the recent study by [37], who did not include DB. Other commonly used functionals could also be included, such as regression model coefficients, non-parametric correlation, and distances between distributions. Furthermore, the study focuses on independent data. Related work on hierarchical, temporal, spatial, and other dependencies is sparse, but mostly in favor of bootstrapping. And in more complex problems bootstrapping is often the only viable approach.
4.1 Practical implications
One general implication of these results for practitioners is that DB is superior to BCa. Considering that the DB is also conceptually simpler and easier to implement than the BCa, it should be a standard feature in every bootstrap library. This is currently not the case. For example, SPSS and popular bootstrap libraries from Python (scipy.stats.bootstrap, bootstrapping) and R (boot) do not implement any form of iterated bootstrap. While the boot package does implement the studentized bootstrap, which involves inner bootstrap sampling, this method performed poorly in our simulation study.
The underutilization of the double bootstrap could be attributed to the misconception that BCa is superior or to the DB being more computationally intensive. Due to its quadratic time complexity, the double bootstrap will not scale to large sample sizes as well as BCa. However, with advances in computation, this is no longer a significant issue. It can also be argued that once a sample size is large enough to make the double bootstrap prohibitive, the choice of method becomes moot, as even the percentile bootstrap will likely perform well enough. Additionally, the double bootstrap is easy to parallelize, and efficient approximations have been developed, such as [53, 54].
In critical applications of statistics, there is no substitute for thoroughly understanding the statistical task at hand, carefully choosing the most appropriate method, and applying it correctly. However, if a single method had to be recommended, the double bootstrap appears to be the best choice. The two scenarios where particular care should be taken are with very small sample sizes and when estimating extreme quantiles, both of which are arguably less common in practice.
The double bootstrap can also play an important role in the teaching of applied statistics. In particular, it can be a valuable tool in a bootstrap-centric introduction to applied statistics for a broad audience of individuals who will use statistics but will only receive one or two courses of formal statistical training. For example, in most empirical sciences.
A basic understanding of standard statistical functionals is required, regardless of what method is used to quantify uncertainty. For example, the mean (average) has to be understood, before bootstrapping, t-distribution-based, or any other approach to constructing CIs can be used. Once functionals are understood, percentile bootstrap is conceptually very simple, can be implemented in a few lines of code, and then applied in the same way to any functional. This is an order of magnitude less complex and time-consuming than teaching a different approach for every functional.
The main disadvantage of the percentile bootstrap is its suboptimal performance in practice. Consequently, a student transitioning into a practitioner must either switch to more complicated bootstrap approaches, such as BCa, or apply parametric approaches. The double bootstrap alleviates this issue, similar to the percentile bootstrap. With just a few more lines of code, we can achieve excellent performance, often surpassing methods that are more complex to understand and implement.
This bootstrap-centric approach is less pedagogically complex than learning several different approaches, and it is computationally not an issue on modern hardware. Lack of software support might be an issue, but a minor one, because percentile and double bootstrap are easy to implement. We agree with [1] that the main reasons that such a shift has not occurred is the lack of teaching materials and legitimate concerns for backwards compatibility. That is, the concern with not teaching the students the methods that are currently dominant statistical practice. This study contributes to alleviating those concerns, showing that bootstrapping, in particular double bootstrap, is overall at least as good as, if not superior to, standard parametric approaches. However, the logical next step is to develop and test an introductory course in applied statistics that relies solely on the bootstrap.
@@ Acknowledgment
This work was supported by the Slovenian Research and Innovation Agency under Grants P2-0442 and J5-60084.